# Nexus: Taming Throughput-Latency Tradeoff in LLM Serving via Efficient GPU Sharing

### v2

@@ Abstract.
Current prefill–decode (PD) disaggregation is typically deployed at the level of entire serving engines [^1] We use the term serving engine to denote a unit of GPUs that manages exactly one complete copy of the model weights., assigning separate GPUs to handle prefill and decode phases. While effective at reducing latency, this approach demands more hardware. To improve GPU utilization, Chunked Prefill mixes prefill and decode requests within the same batch, but introduces phase interference between prefill and decode.
While existing PD disaggregation solutions separate the phases across GPUs, we ask: can the same decoupling be achieved within a single serving engine? The key challenge lies in managing the conflicting resource requirements of prefill and decode when they share the same hardware. In this paper, we first show that chunked-prefill requests cause interference with decode requests due to their distinct requirements for GPU resource. Second, we find that GPU resource exhibits diminishing returns—beyond a saturation point, increasing GPU allocation yields negligible latency improvements. This insight enables us to split a single GPU’s resources and dynamically allocate them to prefill and decode on the fly, effectively disaggregating the two phases within the same GPU.
Across a range of models and workloads, our system Nexus achieves up to 2.2 $\times$ higher throughput, 20 $\times$ lower TTFT, and 2.5 $\times$ lower TBT than vLLM, and outperforms SGLang with up to 2 $\times$ higher throughput, 2 $\times$ lower TTFT, and 1.7 $\times$ lower TBT, and achieves 1.4× higher throughput than vLLM-disaggregation with only half the number of GPUs.
@@ 1. Introduction
Transformer-based (Vaswani et al., 2017) Large Language Models (LLMs) (OpenAI, 2025; DeepSeek, 2025; Brown et al., 2020a; Google, 2025) have achieved state-of-the-art performance on a wide range of tasks, from natural language understanding to code synthesis (OpenAI, 2023; Bai et al., 2023a; Yang et al., 2024, 2025; Team et al., 2025; Guo et al., 2025; Liu et al., 2024a; Guo et al., 2024; Luo et al., 2025). The success has also driven their integration into latency-sensitive applications such as chatbots (OpenAI, 2025; Google, 2025; DeepSeek, 2025; Team et al., 2025; OpenAI, 2023; Guo et al., 2025), search assistants (perplexity, 2025), and AI-augmented IDEs (Cursor, 2025). In interactive settings, even small delays matter a lot: humans perceive latencies above one second as disruptive (Zhong et al., 2024), and sub-second improvements has been shown to substantially boost engagement (GigaSpaces, 2023). Therefore, latency has become a critical performance metric for LLM serving.
LLM inference consists of two distinct stages with heterogeneous resource demands: prefill and decode. In the prefill stage, the model processes the entire prompt in a single forward pass to produce the first output token while populating the key-value (KV) cache. This stage is typically compute-bound (Zhu et al., 2024; Patel et al., 2024), dominated by large matrix multiplications that fill GPU compute units. In contrast, the decode stage generates tokens one at a time, attending to all previously cached KV states. The decode stage, processing one token per request in each forward pass, features lightweight compute but requires reading the full model weights and the entire KV cache, making it heavily constrained by memory bandwidth (Zhu et al., 2024; Patel et al., 2024; Kamath et al., 2025; Kwon et al., 2023).
These two stages naturally give rise to two critical latency metrics in LLM serving: time-to-first-token (TTFT), the delay before the first output token is produced, determined by the prefill stage; and Time-between-Tokens (TBT), the latency between subsequent tokens, dictated by each iteration of the decode loop. These metrics directly impact the user experience in latency sensitive applications. Optimizing TTFT and TBT has driven a wave of innovations in LLM serving systems, spanning scheduling, batching, and kernel design (Agrawal et al., 2024; Zheng et al., 2023; Kwon et al., 2023; Wu et al., 2023; Sun et al., 2024; Ye et al., 2025; Zhong et al., 2024; Nvidia, 2024; Miao et al., 2024a, b; Oh et al., 2024; Mei et al., 2025; Kamath et al., 2025; Yuan et al., 2025; Yu et al., 2025; Feng et al., 2025; Xu et al., 2025).
@@
(b) PD Disaggregated.
(c) Intra-engine disaggregation.
@@
Figure 1. Design evolution of LLM inference systems. Comparison between monolithic, disaggregated, and intra-engine disaggregated designs. Ap is the prefill phase of request A; Bd, Cd, and Dd are the decode phases of requests B, C, and D.
@@
Existing LLM serving systems fall into two broad classes based on how they place prefill and decode execution (Figure 1(a), 1(b)).
@@
Monolithic systems (Kwon et al., 2023; Zheng et al., 2023; Zhu et al., 2024; Agrawal et al., 2024; Yu et al., 2025; Mei et al., 2025; team, 2025) execute both stages within a single engine. To improve utilization, recent designs like Sarathi-Serve (Agrawal et al., 2024) adopt chunked prefill, where long prompts are split into shorter chunks and batched alongside decode tokens. This improves token throughput and reduces TBT by increasing batch efficiency. However, this design mixes compute heavy prefill and memory sensitive decode operations in the same batch, causing interference between the prefill and decode and increasing TBT (see Section 3.1 for details).
@@
Disaggregated systems (Zhong et al., 2024; Patel et al., 2024; Qin et al., 2025; Hu et al., 2024) assign prefill and decode to separate engines, transferring KV cache data between them. This eliminates interference and achieves consistently low TTFT and TBT. But it sacrifices efficiency (Feng et al., 2025): decode engines are often underutilized, and multi-GPU deployment incurs additional hardware and communication overhead, especially when KV state is large.
This paper asks a simple question: Can a single engine achieve low TTFT and TBT without sacrificing GPU utilization? We answer “yes” by introducing intra-engine logical disaggregation, which separates prefill and decode execution within a single serving engine. This design is grounded in three observations. First, mixing compute heavy prefill and memor bound decode in chunked batches creates fine-grained interference that increases TBT. Second, both stages exhibit diminishing returns beyond moderate compute allocations overprovisioning is wasteful, and equal partitioning is inefficient. Third, modern accelerators (Nvidia, 2025c, a) now provide sufficient on device memory and compute to support disaggregation within a single engine, avoiding the communication overheads of cross-GPU designs.
However, enabling this form of disaggregation raises new challenges. Unlike traditional multi-engine systems, we must partition and schedule GPU resources dynamically within a single serving engine, while minimizing contention and adapting to workload shifts in sub-second timescales. Prefill and decode pressure evolve continuously with prompt length, KV cache footprint, and request mix, requiring fine-grained orchestration that scales.
To address these challenges, we present Nexus (Figure 1c), a monolithic LLM serving engine that achieves intra-engine prefill–decode disaggregation. Nexus is built on three components: (1) A lightweight analytical cost model predicts latency as a function of resource allocation, prompt length, and cache usage. (2) A greedy search algorithm consults this model to select low-latency resource partitions in real time, requiring only a few closed-form evaluations per update. (3) Two phase-specific schedulers, shortest-prompt-first for prefill and FCFS for decode, exploit their differing characteristics to optimize TTFT and TBT.
Together, these mechanisms allow Nexus to match the high utilization of monolithic designs while achieving the isolation benefits of disaggregated systems—without incurring cross-device transfers or relying on additional hardware.
@@
We implement Nexus by extending vLLM (Kwon et al., 2023), adding fine-grained resource partitioning and concurrent prefill–decode execution within a single engine. Nexus runs on commodity GPUs without kernel modification, supports decoder-only LLMs, and requires no specialized hardware.
@@
  • We identify the limitations of existing solutions in serving LLMs and propose intra-engine PD disaggregation as a solution.
  • We develop a lightweight, adaptive scheduling mechanism that enables sub-second resource repartitioning and phase-aware prioritization, making intra-engine PD disaggregation practical under real-world, dynamic workloads.
  • We implement Nexus as a drop-in vLLM extension and evaluate it on production-scale LLMs and traffic. Nexus achieves up to 2.2$\times$ and 2$\times$ higher throughput, 20$\times$ and 2$\times$ lower TTFT, and 2.5$\times$ and 1.7$\times$ lower TBT than vLLM and SGLang respectively and outperforms disaggregated vLLM with half the GPU resources.
@@ 2.4. Batching in LLM Serving
Modern LLM serving systems must coordinate prefill and decode phases with fundamentally different latency objectives. Prefill performance directly determines TTFT, while decode responsiveness governs TBT. Since both phases contend for GPU resources. Existing LLM serving systems fall into two broad categories: Monolithic Systems and PD Disaggregation systems, depending on how they manage prefill and decode execution.
Monolithic System
Monolithic system (Kwon et al., 2023; Zheng et al., 2023) such as Sarathi-Serve (Agrawal et al., 2024)(Figure 1(a)) splits long prompts into fixed-size chunks and mixing them with decode tokens in a shared batch queue, these systems construct mixed-phase batches that improve GPU utilization and reduce head-of-line blocking.
PD Disaggregation
To isolate prefill and decode execution, some systems adopt prefill–decode disaggregation (Zhong et al., 2024; Patel et al., 2024; Hu et al., 2024; Qin et al., 2025)(Figure 1(b)), assigning each phase to separate serving engine. This strategy enables independent scheduling and eliminates interference between prefill and decode, particularly in multi-GPU environments.
@@ 3. Motivation
Modern LLM serving must balance two asymmetric phases: compute-heavy prefill and memory-bound decode. In the following section, we examine limitations of current systems that mix or isolate these phases (§3.1), explore how both exhibit diminishing returns with added resources (§3.2), and highlight the inefficiency of static resource partitioning under memory bandwidth contention (§3.3). Together, these insights motivate intra-engine disaggregation with GPU resourcereallocation on the fly.
Setup. Unless stated otherwise, we evaluate with Long Data Collections Workload (§6.1) using Qwen2.5–3B on a single NVIDIA L20 GPU. Requests follow a Poisson arrival at 2.5 req/s. We use NVIDIA MPS (Nvidia, 2025e) to control SM partitioning.
@@ 3.1. Limitation of Existing Solutions
Existing LLM serving systems can fall into two architectural categories: monolithic execution and disaggregation.
Disaggregated systems (Zhong et al., 2024; Qin et al., 2025; Patel et al., 2024) places prefill and decode to separate engines, eliminating resource contention and stabilizing latency. However, the clean separation come with steep hardware costs: multiple engines must maintain the full model replica, prefill engine’s memory is wasted, and decode often underutilize their assigned GPUs. Worse, coordination is non-trivial. Under dynamic workloads, KV cache eviction and recomputation have been shown to severely inflate both TTFT and TBT (Feng et al., 2025).
Monolithic systems (Kwon et al., 2023; Zheng et al., 2023; Agrawal et al., 2024; Yu et al., 2025) adopt chunked prefill, batching prefill and decode requests to improve utilization. Although the strategy increases throughput, it does not account for the distinct compute and memory behavior of each phase, causing phase interference. To quantify this, we categorize batches into prefill-only, decode-only, and mixed, and measure their iteration times (Figure 4(a)). Despite having similar token counts, mixed batches average 250 ms, compared to just 15 ms for decode-only batches.
This slowdown arises from full prefill computation (e.g., KQV projection, FFN) blocking lightweight decode kernels in the same batch. As shown in Figure 4(b), linear kernel latency in mixed batches is up to 10$\times$ higher than in decode-only batches. Since decode cannot proceed until all prefill kernels are completed, this inflates TBT by more than 8$\times$.
{mdframed}
[backgroundcolor=gray!7, linewidth=0.8pt] Insight 1. Chunked prefill improves utilization but introduces fine-grained interference within batches. PD disaggregation avoids this but wastes resources. We aim to achieve both isolation and efficiency within a single engine.

### v3

# Proactive Intra-GPU Disaggregation of Prefill and Decode in LLM Serving

@@ Abstract.
Monolithic serving with chunked prefill improves GPu utilization by batching prefill and decode together but suffers from fine-grained phase interference. Engine-level [^1] We use the term engine to denote a unit of GPUs that manages exactly one complete copy of the model weights. prefill–decode (PD) disaggregation avoids interference but incurs more hardware and high coordination overhead. Prior Intra-GPU disaggregation multiplexs prefill and decode within a single GPU, which uses SLO-based tunning via heuristics derived from offline profiling or feedback loop to allocate GPU resource but respond reactively to performance issues rather than anticipating them, limiting the ability to adapt efficiently under dynamic workloads.
We ask: can we achieve proactive Intra-GPU disaggregation that adapts effectively to dynamic workloads? The key challenge lies in managing the conflicting resource demands of prefill and decode under varying workload conditions. In this paper, we first show that GPU resources exhibit diminishing returns—beyond a certain saturation point, increasing allocation provides minimal latency benefits. Second, we observe that under dynamic workloads, memory bandwidth contention becomes a critical performance bottleneck. These insights motivate a design that splits a single GPU’s resources and dynamically allocates them between prefill and decode phases, while jointly considering compute capacity, memory footprint, and bandwidth contention
Evaluated across diverse LLMs and workloads, our system Nexus achieves up to 2.2$\times$ higher throughput, 20$\times$ lower TTFT, and 2.5$\times$ lower TBT than vLLM; outperforms SGLang by up to 2$\times$; and matches or exceeds the performance of disaggregated vLLM while using only half the GPU resources.
@@ 1. Introduction
(b) Engine-Level PD Disagg.
(c) Intra-GPU PD Disagg.
@@
Figure 1. Design evolution of LLM inference systems. Comparison between monolithic, engine-level disaggregated, and intra-GPU disaggregated designs. Ap is the prefill phase of request A; Bd, Cd, and Dd are the decode phases of requests B, C, and D.
Transformer-based (Vaswani et al., 2017) Large Language Models (LLMs) (OpenAI, 2025; DeepSeek, 2025; Brown et al., 2020a; Google, 2025) have achieved state-of-the-art performance on a wide range of tasks, from natural language understanding to code synthesis (OpenAI, 2023; Bai et al., 2023a; Yang et al., 2024, 2025; Team et al., 2025; Guo et al., 2025; Liu et al., 2024a; Guo et al., 2024; Luo et al., 2025). The success has also driven their integration into latency-sensitive applications such as chatbots (OpenAI, 2025; Google, 2025; DeepSeek, 2025; Team et al., 2025; OpenAI, 2023; Guo et al., 2025), search assistants (perplexity, 2025), and AI-augmented IDEs (Cursor, 2025). In interactive settings, even small delays matter a lot: humans perceive latencies above one second as disruptive (Zhong et al., 2024), and sub-second improvements has been shown to substantially boost engagement (GigaSpaces, 2023). Therefore, latency is a critical performance metric for LLM serving.
LLM inference consists of two distinct stages with heterogeneous resource demands: prefill and decode. In the prefill stage, the model processes the entire prompt in a single forward pass to populate the key-value (KV) cache and produce the first output token. This stage is typically compute-bound (Zhu et al., 2024; Patel et al., 2024), dominated by large matrix multiplications that fill GPU compute units. In contrast, the decode stage generates tokens one at a time, attending to all previously cached KV states. The decode stage features lightweight compute but requires reading the full model weights and the entire KV cache, making it heavily constrained by memory bandwidth (Zhu et al., 2024; Patel et al., 2024; Kamath et al., 2025; Kwon et al., 2023).
These two stages naturally give rise to two critical latency metrics in LLM serving: time-to-first-token (TTFT), the delay before the first output token is produced, determined by prefill; and Time-between-Tokens (TBT), the latency between subsequent tokens, dictated by the decode loop. These metrics directly impact the user experience in latency sensitive applications. Optimizing TTFT and TBT has driven a wave of innovations in LLM serving systems, spanning scheduling, batching, and kernel design (Agrawal et al., 2024; Zheng et al., 2023; Kwon et al., 2023; Wu et al., 2023; Sun et al., 2024; Ye et al., 2025; Zhong et al., 2024; Nvidia, 2024; Miao et al., 2024a, b; Oh et al., 2024; Mei et al., 2025; Kamath et al., 2025; Yuan et al., 2025; Yu et al., 2025; Feng et al., 2025; Xu et al., 2025).
@@
Conventional LLM serving systems fall into two broad classes based on placement of prefill and decode (Figure 1(a), 1(b)).
@@
Monolithic systems (Kwon et al., 2023; Zheng et al., 2023; Zhu et al., 2024; Agrawal et al., 2024; Yu et al., 2025; Mei et al., 2025; team, 2025) execute both stages within a single engine. To improve utilization, recent designs like Sarathi-Serve (Agrawal et al., 2024) adopt chunked prefill, where long prompts are split into shorter chunks and batched alongside decode tokens on the same device. This increases batch efficiency and improves token throughput. However, it mixes compute heavy prefill and memory sensitive decode in the same batch, causing interference (see Section 3.1).
@@
Engine-Level PD Disaggregated systems (Zhong et al., 2024; Patel et al., 2024; Qin et al., 2025; Hu et al., 2024) assign prefill and decode to separate engines, transferring KV cache data between them. This eliminates interference and consistently achieves low TTFT and TBT. But it sacrifices efficiency: decode engines are often underutilized, and multi-GPU deployment incurs additional hardware and communication overhead, especially when KV state is large (Feng et al., 2025; Hong et al., 2025).
Motivated by these limitations and enabled by growing GPU capacity (Nvidia, 2025c, a), recent systems have begun to explore intra-GPU PD disaggregation (Figure 1(c)), running prefill and decode concurrently within a single GPU via spatial or temporal multiplexing (Lin et al., 2025; Cui et al., 2025; Hong et al., 2025). These systems demonstrate the feasibility of intra-GPU disaggregation and show isolating phases within a single device can yield performance benefits.
However, existing solutions guide resource allocation using offline profiling, simple inverse-scaling models, or latency-driven control loops. While these approaches can be effective under stable SLOs or predictable traffic, they lack the adaptability required to respond to changing prompt lengths, KV Cache sizes, and evolving batch compositions. Critically, they fail to jointly model the three core resource pressures: compute saturation, memory bandwidth contention, and runtime memory usage. As a result, these systems adjust resource allocations only after detecting latency violations, making them reactive rather than proactive. This passivity limits their ability to anticipate interference and proactively prevent performance collapse under shifting workloads.
These challenges call for a runtime system that can model such dynamics explicitly, and proactively adapt resource allocation and scheduling within sub-second granularity. We present Nexus, a monolithic LLM serving engine that performs intra-GPU PD disaggregation through model-guided scheduling and lightweight per-batch optimization.
Nexus is built on three components: (1) A contention-aware cost model that predicts latency as a function of compute saturation and memory interference. It is parameterized by a lightweight one-time kernel profiling pass per configuration and requires no workload-specific retraining or online feedback tuning. The model feeds into a dual-objective optimization routine with relaxed latency constraints, solved via fast greedy search to determine per-batch SM partitioning. (2) A buffered asynchronous switching mechanism that mitigates instability during SM reallocation by introducing a hysteresis zone. This avoids oscillation from transient workload shifts and ensures smooth partition transitions. (3) Two phase-specific schedulers: a Shortest-Prompt-First (SPF) policy for prefill to minimize TTFT under skewed prompt lengths, and a simple FCFS policy for decode to preserve fairness and responsiveness.
@@
We implement Nexus by extending vLLM (Kwon et al., 2023) with fine-grained SM partitioning and concurrent prefill–decode execution. Nexus runs on commodity GPUs, supports decoder-only LLMs, and requires no specialized hardware.
@@
  • We identify diminishing compute returns and phase-dependent memory bandwidth contention as key system level bottlenecks in intra-GPU LLM serving.
  • We develop a lightweight, contention-aware runtime framework that models these effects analytically and solves a dual-objective GPU allocation problem via fast greedy search without relying on workload-specific profiling or SLO feedback.
  • We implement Nexus as a drop-in vLLM extension and evaluate it on production-scale LLMs and traffic. Nexus achieves up to 2.2$\times$ and 2$\times$ higher throughput, 20$\times$ and 2$\times$ lower TTFT, and 2.5$\times$ and 1.7$\times$ lower TBT than vLLM and SGLang respectively, and outperforms disaggregated vLLM with half the GPU resources.
@@ 2.4. Batching in LLM Serving
LLM serving systems batch requests to improve throughput by maximizing GPU utilization. However, batching must account for two distinct stages of execution: prefill, which processes the full input prompt and is compute-heavy and batch-friendly; and decode, which generates one token at a time and is memory-bound, latency-sensitive, and harder to parallelize. These phases differ in resource demand and can interfere when co-scheduled on the same GPU, motivating a spectrum of batching and disaggregation strategies.
We categorize prior designs based on how they separate or interleave prefill and decode execution:
Monolithic Systems. Monolithic system (Kwon et al., 2023; Zheng et al., 2023) such as Sarathi-Serve (Agrawal et al., 2024) (Figure 1(a)) splits long prompts into fixed-size chunks and mixing them with decode tokens in a shared batch queue. These systems construct mixed-phase batches that improve GPU utilization and reduce head-of-line blocking.
Engine-Level PD Disaggregation. To isolate prefill and decode execution, some systems adopt engine-level prefill–decode disaggregation (Zhong et al., 2024; Patel et al., 2024; Hu et al., 2024; Qin et al., 2025)(Figure 1(b)), assigning each phase to separate serving engine. This strategy enables independent scheduling and eliminates interference between prefill and decode, particularly in multi-GPU environments.
Intra-GPU PD Disaggregation. A newer class of systems (Lin et al., 2025; Cui et al., 2025; Hong et al., 2025) disaggregates prefill and decode within a single GPU. These systems aim to combine the efficiency of monolithic designs with the phase isolation of disaggregated ones. They typically rely on latency-targeted orchestration, offline profiling, or isolation-based scheduling. Our work follows this line of design but focuses on model-driven, fine-grained proactive scheduling to enable per-batch adaptation.
@@ 2.5. GPU Execution Model
GPU Multiplexing. Modern GPUs support spatial multiplexing via three main mechanisms: NVIDIA MPS, MIG, and CUDA GreenContext (Nvidia, 2025e, d, b). MPS allows concurrent kernel execution across processes with partial SM sharing, while MIG enforces strict hardware-level SM partitioning with fixed memory allocations. GreenContext enables logical SM isolation within a single GPU process by assigning execution streams to non-overlapping SM sets if resource permits. While these mechanisms offer fine-grained compute isolation, they do not virtualize memory bandwidth: all SM partitions share the same L2 cache and DRAM channels. Thus, concurrent execution of kernels can still suffer from contention on the memory subsystem, degrading performance.
@@ 3. Motivation
Modern LLM serving must coordinate two asymmetric phases: compute-heavy prefill (Zhu et al., 2024; Patel et al., 2024) and memory-bound, latency-sensitive decode (Zhu et al., 2024; Patel et al., 2024; Kamath et al., 2025; Kwon et al., 2023). These phases contend for GPU resources in different ways, complicating scheduling and batching decisions.
In this section, we examine the limitations of existing systems that either mix or isolate these phases (§3.1); demonstrate how both prefill and decode exhibit diminishing returns under increased compute allocation (§3.2); and show that static GPU partitioning fails to account for dynamic memory contention that emerges at runtime (§3.3). Together, these observations motivate the need for a runtime system that can adaptively allocate GPU resources while explicitly accounting for compute saturation and memory bandwidth interference.
Setup. Unless otherwise stated, we evaluate using the Long Data Collections workload (§6.1) on Qwen2.5–3B running on a single NVIDIA L20 GPU. Requests arrive via a Poisson process at 2.5 req/s. We use NVIDIA MPS (Nvidia, 2025e) to enforce strict SM partitioning.
@@ 3.1. Limitation of Existing Solutions
Modern LLM serving systems adopt three high-level strategies to manage prefill and decode execution: monolithic batching, engine-level disaggregation, and intra-GPU disaggregation. While each aims to mitigate resource contention between compute-heavy prefill and latency-sensitive decode, they introduce trade-offs that limit flexibility and efficiency.
Monolithic Systems. Monolithic systems (Kwon et al., 2023; Zheng et al., 2023; Agrawal et al., 2024; Yu et al., 2025) batch prefill and decode requests together, often using chunked prefill to improve utilization. However, this co-scheduling introduces interference between resource-asymmetric stages. We categorize batches into prefill-only, decode-only, and mixed, and measure their iteration times (Figure 4(a)). Despite having similar token counts, mixed batches average 250 ms, compared to just 15 ms for decode-only batches.
This slowdown arises from prefill computation (e.g., KQV projection, FFN) blocking lightweight decode kernels within a batch. As shown in Figure 4(b), decode kernel latency increases up to 10$\times$ when co-scheduled with prefill, inflating TBT by more than 8$\times$.
Engine-Level Disaggregation. Engine-level disaggregation (Zhong et al., 2024; Qin et al., 2025; Patel et al., 2024) assigns prefill and decode to separate GPU engines, eliminating phase interference. While this stabilizes latency, it incurs steep hardware and coordination costs: model replication, underutilized decode engines, and KV-cache transfer overhead. These systems also assume fixed phase placement and lack runtime adaptation. Under bursty or decode-heavy conditions, coordination failures can trigger cache eviction, recomputation, or out-of-memory (OOM) errors (§6.2.2(Feng et al., 2025; Hong et al., 2025).
Intra-GPU Disaggregation. Recent systems such as Bullet (Lin et al., 2025), Drift (Cui et al., 2025), and semi-PD (Hong et al., 2025) disaggregate prefill and decode execution within a single GPU, reducing hardware cost and avoiding cross-GPU KV transfers. These designs demonstrate that intra-GPU disaggregation is feasible, but often rely on simplified assumptions. Bullet uses SLO-guided feedback to adjust resource allocations based on profile-derived latency estimators. Drift assumes contention-free execution and applies gang scheduling with phase tagging. Semi-PD fits inverse-scaling latency models and applies windowed feedback to guide SM partitioning. While effective under predictable workloads, these methods are inherently reactive and do not explicitly model two critical system-level factors: (1) diminishing returns in compute allocation, and (2) dynamic memory bandwidth contention. As we show in the next sections, both effects vary significantly across phases and should be anticipated to enable robust, fine-grained adaptation.